OpenStudy (anonymous):

Solve for x: −3|x + 7| = −12

OpenStudy (anonymous):

Let's solve your equation step-by-step. −3(|x+7|)=−12 Step 1: Divide both sides by -3. −3(|x+7|) −3 = −12 −3 |x+7|=4 Step 2: Solve Absolute Value. |x+7|=4 We know either x+7=4 or x+7=−4

OpenStudy (dan815):

-3*WHAT = -12

OpenStudy (anonymous):

do you know the next step?

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

yess i am still here @ROMNAY

OpenStudy (anonymous):

now just simply solver for x with those last two equations and then you get your asnwers